# GALLa: Graph Aligned Large Language Models for Improved Source Code Understanding

### v1

@@ 4.4 Ablation Studies
As a control experiment, we additionally trained a model on the mixture of downstream tasks and the source code from the graph alignment data, but did not provide it with the graph information - in other words, Graph2Code with only the code but not the graph, as illustrated in Figure 3, which is similar in essence to continual pretraining on the source code data. The results are given in the last row in Table 5, which suggests that the graph information is indeed helping the LLM to better understand programs.
To show that the proposed GALLa framework can be applied to various GNN encoders and adapter modules, we also conducted another two sets of experiments in the MFT setting with Phi-1: based on the settings in the main experiments, we 1) replaced the cross-attention adapter with a 3-layer MLP (which has a similar parameter count to the cross-attention layer), and 2) replaced the graph encoder with a MagNet (Zhang et al., 2021). For these two experiments, we repeated the complete two stages of GALLa training, and the results for using an MLP as the adapter are presented in Table 6. Compared with the results using cross-attention from Table 3, we surprisingly find that the MLP adapter leads to a slightly better performance, which is counter-intuitive as MLP does not allow for information exchange between the nodes. We hypothesize that this is because with the same parameter count, MLP has more layers than cross-attention (3 vs. 1 in our case), thus having a stronger expressivity.
The results of using a different graph encoder - MagNet (Zhang et al., 2021) - when finetuning Phi-1 with GALLa are presented in Table 7. Compared with the results using DUPLEX as the graph encoder in Table 3 in the main paper, we observe that MagNet can similarly improve the LLM’s performance on all five downstream tasks. Together with the experiments of using different LLMs (Table 3) and different adapters (Table 6), this confirms that GALLa is a flexible framework where each of the three modules - GNN encoder, adapter, and LLM decoder - can be replaced upon demand.
@@ A.3 Universal AST
Typically, ASTs are specific to the programming language in question. In other words, different programming languages would have different types of nodes in their ASTs - for example, a for loop node in Python may have different semantics from a for loop in C++. However, in the context of our work, such language-specific specifications may be detrimental to the cross-language alignment or generalization capabilities of the models.
Thus, in this work we used a special type of AST (and DFG, as they share the same set of nodes): Universal AST (UAST). In UAST, the specifications of all node types are designed to be as language-independent as possible. For example, for loops from different languages are not distinguished from each other, and only the most basic semantics of a for loop is retained in this abstract node type.
@@
Appendix B Statistical Tests
The complete list of p values of the statistical differences between the main results and the baselines is given in Table 10.
Model Setting Trans (Ja2Py/Py2Ja) Clone Repair Sum (Java/Py/JS) Defect CodeGen 350M G2C 0.0862/0.0000 0.0000 0.0000 0.0114/0.0000/0.3684 0.0000 G2C + GQA 0.0000/0.0001 0.0000 0.0000 0.0111/0.0011/0.4332 0.0000 StarCoder 1B G2C 0.0057/0.0140 0.0000 0.0000 0.0000/0.0000/0.0000 0.0000 G2C + GQA 0.0057/0.0140 0.0000 0.0000 0.0000/0.0000/0.0000 0.0000 Phi-1 1.3B G2C 0.0000/0.0000 0.0000 0.0000 0.0090/0.0595/0.5485 0.0000 G2C + GQA 0.0009/0.0000 0.0000 0.0000 0.0058/0.0048/0.7818 0.0000 LLaMA3 8B G2C 0.0003/0.0000 0.0000 0.0000 0.5029/0.0000/0.0014 0.0000 G2C + GQA 1.0000/0.0000 0.0000 0.0000 0.0019/0.0147/0.0001 0.0000

### v2

@@ 4.4 Ablation Studies
Contribution of the Graph Information
Observing the previous results, one may raise the question: are these improvements indeed attributable to the alignment of code representations to graphs, or are they simply a result of additional computation expense during the finetuning of LLMs?
To answer this question, we conducted a control experiment where we additionally trained a model on the mixture of downstream tasks and the source code from the graph alignment data, but did not provide it with the graph information - in other words, Graph2Code with only the code but not the graph, as illustrated in Figure 3, which is similar in essence to continual pretraining on the source code data. The results are given in the last row in Table 5, which suggests that the graph information is indeed helping the LLM to better understand programs.
Different Adapters and Graph Encoders
To show that the proposed GALLa framework can be applied to various GNN encoders and adapter modules, we also conducted another two sets of experiments in the MFT setting with Phi-1: based on the settings in the main experiments, we 1) replaced the cross-attention adapter with a 3-layer MLP (which has a similar parameter count to the cross-attention layer), and 2) replaced the graph encoder with MagNet (Zhang et al., 2021). For these two experiments, we repeated the complete two stages of GALLa training, and the results are presented in Table 6.
Compared with the results using cross-attention from Table 3, we surprisingly find that the MLP adapter leads to a slightly better performance, which is counter-intuitive as MLP does not allow for information exchange between the nodes. We hypothesize that this is because with the same parameter count, MLP has more layers than cross-attention (3 vs. 1 in our case), thus having a stronger expressivity.
Similarly, when using MagNet (which is convolution-based) instead of DUPLEX (which is self-attention-based) for finetuning Phi-1, the model’s performance on downstream tasks is also slightly improved. These results of using different adapters and graph encoders, together with the experiments of using different LLMs (Table 3), confirm that GALLa is indeed a flexible framework where each of the three modules - GNN encoder, adapter, and LLM decoder - can be replaced as more advanced models are proposed in the future, leading to better performance on downstream tasks.
@@ 5 Conclusion
Limitations
As we pointed out in the related work section, all existing methods that attempt to combine structural graphs with code language models focus on small, BERT-style models. Since our work is the first to utilize structural graphs to enhance decoder-only LLMs in understanding source code, we are unable to find any existing methods for fair comparison, and thus can only fine-tune the LLMs directly on source code as a baseline. However, our work can serve as a baseline for any future research in graph-enhanced (or more generally, structure-aware) code LLMs.
Due to our limited computational budget, we mostly focused on LLMs around 1B-10B parameters (especially in the ablation studies) and only verified the effectiveness of GALLa on one 14B model. We leave the exploration of even larger structure-aware code LLMs to future work. Also due to the budget constraint, we mainly focused on multi-task training in our main experiments and did not conduct individual hyper-parameter searches for the five downstream tasks. However, we maintained a fair comparison between the baseline and the proposed methods.
Ethical Statements
All the datasets and models used in this paper are publicly available resources. We use them only for research purpose.
Acknowledgments
This paper is supported by Ant Group and the General Program of National Natural Science Foundation of China (62176153).
@@ A.3 Universal AST
Typically, ASTs are specific to the programming language in question. In other words, different programming languages would have different types of nodes in their ASTs - for example, a for loop node in Python may have different semantics from a for loop in C++. However, in the context of our work, such language-specific specifications may be detrimental to the cross-language alignment (e.g. the code translation task) or generalization capabilities (i.e. generalizing to languages not present in the training data) of the models.
Thus, in this work we used a special type of AST (and DFG, as they share the same set of nodes): Universal AST (UAST). In UAST, the specifications of all node types are designed to be as language-independent as possible. Taking the nodes in Table 7 as an example, UAST abstracts language-specific node types into the most basic concepts shared by most programming languages, such as variable, binary expression, and return statement. In total, there are 43 node types in our graph data.
@@
Appendix B GraphQA Prompts
In our main experiments, we used three types of questions for GraphQA: edge prediction, parent prediction, and child prediction. For each type of question, we wrote about ten question templates and ten answer templates, as shown in Figure 4 to 6. When constructing the data, one question template and one answer template are randomly chosen for each sample, as diverse question templates have been found to improve cross-task generalization capability in instruction finetuning (Sanh et al., 2022). The placeholders in the templates are replaced with the actual node contents and types, and the instantiated prompt is then randomly placed before or after the graph tokens.
In our preliminary experiments, we found edge prediction data constructed from AST to bring no improvement, as connected nodes in the AST often have high textual overlap and can be trivially predicted. Thus we only use DFG data for the edge prediction task, while both AST and DFT data are used for the other two tasks. In early trials, we also experimented with and eventually discarded several other types of questions, including node classification (which can be easily done by only looking at the node’s source code) and some tasks that involve counting, such as counting the number of nodes in the graph, the number of edges in the graph, and the number of node types in the graph, which prove to be too difficult for the model.
Appendix C Statistical Tests
To verify the statistical significance of the main results, we conducted Chi squared tests on the four tasks with discrete performance metrics (i.e. code translation, code repair, clone detection, and defect detection) and Wilcoxon signed-rank test on the task with continuous performance metrics (i.e. code summarization). The results are presented in Table 9. Most of the differences are significant, except for CodeGen and Phi-1 on JavaScript summarization, and LLaMA3 on Java-to-Python translation and Python summarization.
Model Setting Trans (Ja2Py/Py2Ja) Clone Repair Sum (Java/Py/JS) Defect CodeGen 350M G2C 0.0862/0.0000 0.0000 0.0000 0.0114/0.0000/0.3684 0.0000 G2C + GQA 0.0000/0.0001 0.0000 0.0000 0.0111/0.0011/0.4332 0.0000 StarCoder 1B G2C 0.0057/0.0140 0.0000 0.0000 0.0000/0.0000/0.0000 0.0000 G2C + GQA 0.0057/0.0140 0.0000 0.0000 0.0000/0.0000/0.0000 0.0000 Phi-1 1.3B G2C 0.0000/0.0000 0.0000 0.0000 0.0090/0.0595/0.5485 0.0000 G2C + GQA 0.0009/0.0000 0.0000 0.0000 0.0058/0.0048/0.7818 0.0000 Qwen2.5-Coder 1.5B G2C 0.0000/0.0000 0.0000 0.0000 0.2176/0.0001/0.0000 0.0000 G2C + GQA 0.0286/0.0000 0.0000 0.0000 0.0008/0.0003/0.0001 0.0000 LLaMA2 7B G2C 0.0037/0.0000 0.0000 0.0000 0.0195/0.0755/0.5454 0.0000 G2C + GQA 0.0000/0.0000 0.0000 0.0000 0.0468/0.1406/0.0147 0.0000 LLaMA3 8B G2C 0.0003/0.0000 0.0000 0.0000 0.5029/0.0000/0.0014 0.0000 G2C + GQA 1.0000/0.0000 0.0000 0.0000 0.0019/0.0147/0.0001 0.0000 Qwen2.5-Coder 14B G2C 0.0000/0.0000 0.0000 0.0000 0.9403/0.0213/0.0000 0.0000 G2C + GQA 0.0509/0.0000 0.0000 0.0000 0.2508/0.6213/0.0000 0.0000
Appendix D Example Outputs with Single-Task Finetuning
In Figure 7, we provide an example of model hallucination after training on only one downstream task in the GALLa framework. The model produces an answer template learned from the GraphQA task instead of responding to the actual question. This issue is mitigated in the multitask finetuning setting.